particles p and Q leave a fixed point A at the same time and travel in the same straight line. the velocity of P after t seconds is 6t(t-3)ms-1 and the velocity of Q after t seconds is (10-2t)ms-1.
1) sketch,on the same axes, velocity-time graph for P and Q for 0<=t<=5
2)verify that P and Q meet after 5 seconds.
3)find the greatest distance between P and Q for 0<=t<=5
1
Expert's answer
2020-04-15T10:16:06-0400

1)



2) After 5 seconds:


"S_1=\int v(t)dt=\int(6t^2-18t)dt=(2t^3-9t^2)|_0^5=25"


"S_2==\int v(t)dt=\int(10-2t)dt=(10t-t^2)|_0^5=25"

3)



"f(t)=v_2t-v_1t"


"f(t)=10t-2t^2-6t^3+18t^2"


"f^\/(t)=-18t^2+32t+10"

"t \approx1"

"f(1)=20"